Question 2 (30 marks)

At 30 June 2018, Spencer Ltd had the following temporary differences:

Asset or liability

Carrying amount ($000)

Tax base ($000)

Computers at cost

300

300

Accumulated depreciation

(60)

(100)

Computers (net)

  240

  200

Accounts receivable

100

100

Allowance for doubtful debts

  (10)

0

Accounts receivable (net)

    90

100

Provision for warranty costs

30

0

Provision for employee benefits (LSL)

20

0

The following information is available for the year ending 30 June 2019.

Statement of comprehensive income for Specialist Engineering Ltd for the year ended 30 June 2019

$000

Revenue

4 000

Cost of goods sold expense

(1 800)

Depreciation expense

(60)

Warranty expense

(90)

Bad and doubtful debts expense

(25)

Other expenses

(1 375)

Profit before tax

650

Additional information:

  • Spencer Ltd depreciates computers over 5 years under accrual accounting but the ATO does it over 3 years. Both the company and the ATO uses the straight-line method.
  • During the year Spencer Ltd wrote off bad debts amounting to $15 000.
  • Warranty costs of $70 000 were paid during the year.
  • No amounts were paid for long-service leave during the year.
  • There was no acquisition of plant and equipment during the year.
  • The tax rate as at 30 June 2018 and 30 June 2019 was 30 per cent.

The following information is extracted from the statement of financial position at 30 June 2019:

$000

Assets

Accounts receivable

120

Allowance for doubtful debts

(20)

Liabilities

Provision for warranty costs

50

Provision for employee benefits (LSL)

30

Required:

  1. Calculate the amount of each of Spencer Ltd’s temporary differences at 30 June 2018, and state whether it is deductible or taxable. (10 marks)
  2. Determine the balance of the deferred tax liability and deferred tax asset as at

30 June 2018 and provide the journal entry. (2 marks)

  1. Calculate Spencer Ltd’s taxable income for the year ended 30 June 2019 and provide the journal entry. (5 marks)
  2. Calculate the deferred tax for the year ended 30 June 2019 and provide the journal entry. (13 marks)
Answer #1

Computer=240-200,sccounts receivable =100-90=10 ,provision for warranty=0-30=-30 carry amount more than tax base it is deferred tax liability and carry amount less than tax base it is Dylan details as follows : computer=0, accumulated depreciation =-160 Dylan and computers 40(dtl), providing for employee benefits -20ie dry, provision for warranty costs =30ie dry accouPnts receivable 10DTA .Balance of deferred tax liability=-160+90=70 and balance of Deferred tax asset =10-10=Nil